OpenStudy (anonymous):

What is the exact volume of the sphere? A. 2481.156 cm3 B. 7263.39 cm3 C. 7790.83 cm3 D. 22807.05 cm3

OpenStudy (anonymous):

u can find it out by formula that is \[\frac{ 4 }{ 3 }\pi r ^{3}\]

OpenStudy (anonymous):

@undeadknight26 @dan815 @peterg2001 @Demonx341 @Legends @jordanjamesbay

OpenStudy (mrnood):

well = since pi is irrational you cannot calculate it 'exactly' the formula is vol = pi r^3 * 4/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r is the radius of sphere

OpenStudy (anonymous):

\[vol=\frac{ 4 }{ 3 }\pi 12.3^{3}\] can u solve it now

OpenStudy (mrnood):

I suppose you could state it 'exactly' by leaving pi in the answer, since it does represent a 'number'

OpenStudy (anonymous):

ok thx guys

OpenStudy (mrnood):

The answer above has used pi = 3.14 (not exactly exact in my book)